OpenStudy (anonymous):

Wyatt purchased a living room set for $4,109 with an 18-month deferred payment plan. The interest rate on the plan was 27.41%. The living room set must be paid off within three years after the deferment period. What is the total cost of the living room set if she made no payments during the deferment period and made only minimum payments after the deferment period? $4,109.00 $6,170.03 $9,117.00 $13,675.50

OpenStudy (anonymous):

@agent0smith this one?

OpenStudy (agent0smith):

I think do it the same way, just with the new numbers. Reuse the formula except P=4109 r=0.2741 n=12 t=3 (3 years) find A. Then divide by 36 (36 months)

OpenStudy (anonymous):

How do I find A?

OpenStudy (agent0smith):

plug those numbers in there\[\Large A = P(1+r/n)^{nt}\]

OpenStudy (agent0smith):

\[\Large A = 4109(1+0.2741/12)^{12*3}\]

OpenStudy (agent0smith):

@brittneyy18rm

OpenStudy (anonymous):

I'm here sorry I was eating haha

OpenStudy (anonymous):

I got 4460.74 :(

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=4109%281%2B0.2741%2F12%29%5E%7B12*3%7D then divide that by 36

OpenStudy (anonymous):

Now I got 257 :(

OpenStudy (agent0smith):

oh wait nm we want the total cost. that's it in the link above

OpenStudy (anonymous):

agh I just want to be lead to the answer :(

OpenStudy (agent0smith):

The steps are at the top, just plug in all the numbers :)

OpenStudy (anonymous):

I plugged it all in and got 9264.86 :((

OpenStudy (agent0smith):

yeah that's close enough to the 9117 answer :) the formula is probably missing something cos i don't know the financial stuff, but it's close enough and i'm pretty sure.

OpenStudy (anonymous):

thank you so much for sticking through this with me

OpenStudy (agent0smith):

:)

OpenStudy (anonymous):

a=4109(1+0.2741/12)^12*3 a=4109(1.022)^12*3 a=4109(1.022)^36 a=4109(3679) then you multiply those and divide by 36 a= 4199, close to A